### v1

# High-dimensional dynamics in low-dimensional networks

Abstract
Many networks that arise in nature and applications are effectively low-dimensional in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks might also be low-dimensional. Indeed, recent results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or noise. However, networks in nature are rarely isolated. We show that recurrent networks with low-rank structure often produce high-dimensional dynamics in the presence of high-dimensional perturbations. Counter to intuition, dynamics in these networks are suppressed in directions that are aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” Our results clarify important, but counterintuitive relationships between a network’s connectivity structure and the structure of the dynamics it generates.
Introduction
Recent work shows that many networks arising in nature and in applications are effectively low-rank in the sense that their connectivity matrices have a small number of larger singular values, while the remaining singular values are much smaller [1]. What are the implications of low-rank network structure on the dimensionality of network dynamics?
Several studies have shown that networks with low-rank structure tend to produce low-dimensional dynamics [1, 2, 3, 4]. However, these studies rely on specific assumptions about the network model. Some studies assume that the low-rank part of the network is weak [2, 4], others assume that the network is self-contained in the sense that it operates in the absence of internal noise or external perturbations or inputs [5, 1], while other work assumes that external perturbations are perfectly aligned with the network’s low-rank structure [3]. In reality, networks in nature are rarely isolated from perturbations, and these perturbations are not necessarily aligned with the network’s low-rank structure.
We use mathematical analysis and computer simulations to show that, in the presence of high-dimensional inputs or perturbations, many networks with low-dimensional structure produce high-dimensional dynamics. Moreover, perhaps counterintuitively, these networks suppress variability in directions that are aligned with their low-rank structure relative to random directions, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity structure promote low-rank suppression and high-dimensional dynamics. Finally, we demonstrate our conclusions in a real world epidemiological network with low-rank structure.
Our conclusions have important implications for the interpretation of low-dimensional structure in recurrent networks. In neuroscience, our results offer a mathematical explanation of the widely observed balance between excitatory and inhibitory synaptic currents [6, 7, 8, 9, 10, 11, 12] and also explain why neural populations generate high-dimensional responses to high-dimensional stimuli [13]. Beyond neuroscience, our results imply that randomly structured perturbations are often most effective at driving activity in networks with low-rank structure.
Results
High-dimensional dynamics and low-rank suppression in a network with rank-one structure
For illustrative purposes, we begin with a simple, linear model (Figure 1a)
$\tau\frac{d{\bf{z}}}{dt}$ $=-{\bf{z}}+W{\bf{z}}+{\bf{x}}.$ (1)
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and we take $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
$W=W_{0}+W_{1}$
where
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
is a rank-one matrix and
$W_{1}=\frac{\rho}{\sqrt{N}}Z$
is a full rank random matrix. Specifically, ${\bf{u}}$ is a random vector with $\|{\bf{u}}\|=1$, $Z$ is an $N\times N$ matrix with entries drawn i.i.d. from a standard normal distribution, and $\rho>0$. If
$|c|\gg\rho$
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are bounded by $2\rho$ (Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane [14, 15]. Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [16]. Since low-rank structure requires $|c|\gg\rho$, stability requires $c<0$ when $\rho\sim\mathcal{O}(1)$. Here, we take $\rho=0.5$ and $c=-10$. We consider more general network structures later.
One might expect a network with connectivity $W$ to produce low-dimensional dynamics due to its low-rank structure. Indeed, a feedforward network with the same connectivity matrix produces approximately one-dimensional dynamics because it amplifies inputs aligned to ${\bf{u}}$ relative to other directions (Supplementary Figure S.1). We next demonstrate that this seemingly natural conclusion does not carry over to recurrent network dynamics.
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, ${\bf{z}}$, which receive external perturbations or input, ${\bf{x}}$. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by ${\bf{u}}$ and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along ${\bf{u}}$. g) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of ${\bf{u}}$.
We first simulated the network with two different input patterns. The “aligned” input pattern was exactly aligned with the low-rank component of $W$,
${\bf{x}}_{\textrm{aligned}}={\bf{u}},$
while the “random” input pattern is aligned with a random direction,
${\bf{x}}_{\textrm{random}}={\bf{u}}_{\textrm{rand}},$
which is generated identically to, but independently from ${\bf{u}}$. For large $N$, ${\bf{u}}$ and ${\bf{u}}_{\textrm{rand}}$ are approximately orthogonal. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network. Perhaps surprisingly, we observed exactly the opposite (Figure 1c): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
To demonstrate some consequences of this effect, we next provided a high-dimensional input, ${\bf{x}}(t)$, that varied smoothly in time. Specifically, each ${\bf{x}}_{j}(t)$ was an i.i.d. realization of a smooth, unbiased, stationary Gaussian process. This can model internal noise or irregular external input to the network.
Conventional wisdom and intuition might lead us to expect the network response to be low-dimensional and dominated by variability in the direction of ${\bf{u}}$, owing to the low-rank structure of $W$. In contrast, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$. The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d). Perhaps surprisingly, this weakest principal component direction was closely aligned to ${\bf{u}}$ (the angle between ${\bf{u}}$ and the corresponding principal component direction was less than $8^{\circ}$). Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of ${\bf{u}}$ was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
We propose the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to dominant the low-rank structure of a network are suppressed by the network’s dynamics. In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional perturbations, as in Figure 1d. In this manuscript, we derive conditions under which low-rank suppression and high-dimensional dynamics arise in low-rank networks. We then show that these conditions are met by many natural low-rank network structures.
The results in Figure 1 contrast with previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In some of these studies, external input or perturbations to the network are assumed to be absent [1, 5] or perfectly aligned with the low-rank component of the network connectivity [3]. In other studies [2, 4], the network is assumed to have only a weak low-rank component ($|c|\sim\rho$). None of these assumptions are consistent with the model considered in Figure 1. See Supplementary Text S.1 for a more detailed review of this previous work and its relation to ours.
A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady state solution to Eq. (1),
${\bf{z}}=[I-W]^{-1}{\bf{x}}.$ (2)
Convergence to the steady state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $[I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near ${\bf{u}}$. Hence, the recurrent network suppresses inputs in the direction of ${\bf{u}}$. A related explanation is that the Jacobian matrix $J=W-I$ of Eq. (1) has a large, negative eigenvalue near $c-1<0$ with associated eigenvector near ${\bf{u}}$ so the dynamics in Eq. (1) are highly compressive along ${\bf{u}}$. These explanations depend on specific properties of the model considered here, but we later describe more general conditions under which low-rank suppression and high-dimensional dynamics occur.
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with network interactions in the direction of ${\bf{u}}$. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of local inputs to each node whereas ${\bf{x}}(t)$ is external input, and ${\bf{z}}(t)$ is the network response. In the qausi-steady state, ${\bf{z}}\approx W{\bf{z}}+{\bf{x}}$. Under low-rank suppression, local input cancels nearly perfectly with external input ($W{\bf{z}}\approx-{\bf{x}}$) in the direction of ${\bf{u}}$ so that the response, ${\bf{z}}$, is weak in this direction (Figure 1g). This cancellation is a key feature of low-rank suppression.
Conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple example of a network with rank-one structure. In Supplementary Text S.2, we derive conditions for low-rank suppression and high-dimensional dynamics in a general class of models that includes all dynamics of the form $\tau d{\bf{z}}/dt={\bf{F}}({\bf{z}},{\bf{x}})$. Here, we summarize and interpret those results.
Low-rank suppression (as in Figure 1c) is a more general phenomenon than high-dimensional dynamics (as in Figure 1d). In Supplementary Text S.2, we show that low-rank suppression arises under two assumptions:
  1. 1.
    Assumptions on singular values. First, we assume that the network is strongly low-rank in the sense that the connectivity matrix $W$ has a small number of large singular values and the rest are much smaller (Figure 2a). Specifically, $W=W_{0}+W_{1}$ where $W_{0}$ has rank $r\ll N$ with large singular values ($\sigma_{W_{0}}\gg 1$) while $W_{1}$ has rank $N$ with small or moderate singular values ($\sigma_{W_{1}}\leq\mathcal{O}(1)$). “Weakly low-rank” networks for which $\sigma_{k}\sim\mathcal{O}(1)$ are studied in other work [2, 17, 4, 18, 19, 20] and do not necessarily produce low-rank suppression (see Supplementary Text S.1).
  2. 2.
    Assumptions on dynamics. In addition, our analysis relies on a linearization of the dynamics around a stable equilibrium. This requires that ${\bf{z}}(t)$ settles to a stable equilibrium whenever ${\bf{x}}(t)=0$, and that ${\bf{x}}(t)$ is sufficiently weak or the dynamics of ${\bf{z}}(t)$ are sufficiently linear. Nonlinear dynamics away from a stable equilibrium can produce any dynamics in principle [21, 22, 23] (see Supplementary Figures S.7 and S.8 for examples of unstable network states).
Under these two assumptions, low-rank suppression emerges in the sense that there are inputs, ${\bf{x}}$, approximately parallel to the column space of $U$, that produce much weaker responses than inputs at random directions (as in Figure 1c). The emergence of high-dimensional dynamics (as in Figure 1d) requires additional assumptions:
  1. 3.
    Assumptions on input perturbations. We assume that ${\bf{x}}(t)$ is a smooth, high-dimensional stochastic input. In simulations, we take each ${\bf{x}}_{j}(t)$ to be an i.i.d. Gaussian process. Low-dimensional inputs, ${\bf{x}}(t)$, can produce low-dimensional dynamics in networks that would otherwise produce high-dimensional dynamics (Supplementary Figure S.9).
  2. 4.
    Conditions on the alignment of singular vectors. Under the assumptions above, the emergence of high-dimensional dynamics depends on the structure of the low-rank part, $W_{0}$, of the connectivity. Specifically, let
    $W_{0}=U\Sigma V^{T}$
    be the singular value decomposition of $W$ with $U$ and $V$ being $N\times r$ orthonormal matrices. We define the $r\times r$ recurrent alignment matrix [3, 24],
    $P=U^{T}V.$
    which measures the alignment between the left and right singular vectors of $W_{0}$, i.e., the columns of $U$ and $V$. High-dimensional dynamics can only emerge when no singular value of $P$ is small, i.e., when $\sigma_{P}=\mathcal{O}(1)$ for all singular values of $P$.
To better understand this condition on $P$, we consider simulations of a model that is widely used in computational neuroscience to model a recurrently interacting network of neurons [16, 25, 2, 26],
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (3)
We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 2a) and we fix
$U=\left[\begin{array}[]{cc}{\bf{u}}_{1}&{\bf{u}}_{2}\end{array}\right]$
while exploring different choices of $V$.
We first consider the special case that $W_{0}$ is normal, so ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$. In this case, $P$ is diagonal with $|P_{kk}|=1$ (Figure 2b) so all singular values of $P$ are $\sigma_{P}=1$ and the network exhibits high-dimensional dynamics (Figure 2c). Since $W_{0}$ is normal, its eigenvalues are real and equal in magnitude to the singular values, $|\lambda_{W_{0}}|=\sigma_{W_{0}}$. Since $\sigma_{W_{0}}\gg 1$, stability of the dynamics requires that $\lambda_{W_{0}}=-\sigma_{W_{0}}$ and therefore
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&-{\bf{u}}_{2}\end{array}\right].$
Dynamics on the column space of $U$ are strongly suppressed due to the large, negative eigenvalues.
Figure 2: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal. c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics. d,e) Same as b,c except $W_{0}$ is EP, but not normal. f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP. h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq. (4) with $c=0.5$).
However, low-rank suppression does not require that $W_{0}$ is normal (${\bf{v}}_{k}=\pm{\bf{u}}_{k}$). Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [27]. All normal matrices are EP, but an EP matrix is not necessarily normal. An example of a non-normal, EP matrix with rank 2 is given by
$V$ $=\left[\begin{array}[]{cc}{\bf{u}}_{2}&-{\bf{u}}_{1}\end{array}\right].$
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 2d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$. Regardless, all singular values of $P$ are $\sigma_{P}=1$, and therefore the network produces low-rank suppression and high-dimensional dynamics (Figure 2e).
To obtain an example of a network that produces low-dimensional dynamics, we take
$V$ $=\left[\begin{array}[]{cc}{\bf{v}}_{\perp}&-{\bf{u}}_{2}\end{array}\right].$
where ${\bf{v}}_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$. In this case, $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 2f). Because of the singular value at zero, the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 2g).
Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap, for example
$V$ $=\left[\begin{array}[]{cc}\sqrt{1-c}{\bf{v}}_{\perp}-\sqrt{c}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (4)
where $10<c<$ 1. In this case, $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and another at $\sigma_{P}=1$ (Figure 2h). As long as $c$ is not asymptotically close to zero, the network produces low-rank suppression and high-dimensional dynamics (Figure 2i).
In summary, the class of effectively low-rank matrices, $W$, that produce low-dimensional dynamics in the presence of high-dimensional inputs is the very specific class of matrices for which the alignment matrix, $P=U^{T}V$, of the low-rank part is singular or nearly singular in the sense that it has a small or zero singular value.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.2). However, many networks arising in nature and applications do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
Biased weights and modular networks
So far, we considered example networks unbiased weights, $E[W_{jk}]=0$, which is common in many modeling studies, but many networks in nature have weights with non-zero mean. Biased weights can produce low-rank structure. As a simple example, consider a random network with independent weights satisfying
$E[W_{jk}]=m\;\textrm{ and }\;\textrm{std}(W_{jk})=s.$
If $m$ and $s$ scale similarly to each other, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}\sim\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure because $\sigma_{1}/\sigma_{2}\sim\mathcal{O}(\sqrt{N})$. The dominant rank-one part of such a matrix has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.3).
Figure 3: Low-rank suppression and excitatory-inhibitory balance in a modular network. a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure. c) The alignment matrix shows that the network is EP, but not normal. d) An input that is constant within each block (red) is aligned to the low-rank part, but a random input (gray) is not. e) Response magnitude is suppressed for the aligned input relative to the random input. f) The excitatory (positive; red) component of an input balances with the inhibitory (negative; blue) component to produce a much smaller total (gray) component, a widely observed phenomenon in neural circuits.
More generally, modular structure arises in many natural settings [1, 28]. Specifically, many networks in nature represent the interaction between $n$ populations, and mean connection weights between populations are often non-zero. The adjacency matrices of modular networks can be arranged to have a block structure
$W=\left[\begin{array}[]{ccc}W^{1,1}&\cdots&W^{1,n}\\ W^{2,1}&\cdots&W^{2,n}\\ \vdots&\cdots&\vdots\\ W^{n,1}&\cdots&W^{n,n}\end{array}\right]$
where $W^{a,b}$ is a sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
$E[W_{jk}]=m_{ab}\;\textrm{ and }\;\textrm{std}(W_{jk})=s_{ab}.$
If we assume that each population has $\mathcal{O}(N)$ members and that $m_{ab}$ scale similarly to $s_{ab}$ for large $N$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(\sigma_{ab}\sqrt{N})$. Hence, large modular networks with biased weights naturally produce low-rank structure [1].
Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is the element-wise expectation of $W_{0}$, which is constant within each block, so $W_{0}$ has rank at most $n$. In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the $n$ indicator vectors of the $n$ populations. Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics in response to high-dimensional perturbations.
As a specific example of a modular network, we consider a model of a local neuronal network in the cerebral cortex. Cortical neurons obey Dale’s Law: All outgoing connection weights from a particular neuron have the same sign, positive for excitatory neurons and negative for inhibitory neurons, and mean connection weights also depend on the postsynaptic neuron type [29, 30]. These properties produce a modular structure in which the columns of the adjacency matrix corresponding to excitatory neurons are non-negative, while the columns corresponding to inhibitory neurons are non-positive (Figure 3a). We consider the case in which $m_{ab}$ and $s_{ab}$ scale like $1/\sqrt{N}$, consistent with experiments [31] and theoretical work [32, 33, 34, 35, 3, 36].
This network has an effective rank-two structure (Figure 3b): Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$ and the low-rank part of the connectivity matrix is EP (Figure 3c), so our theory predicts that the network exhibits low-rank suppression and high-dimensional responses to high-dimensional inputs. The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
$\textrm{col}(U)=\textrm{col}(V)=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}.$
Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 3d,e) consistent with recordings from monkey motor cortex [36] and previous theoretical work on balanced network models [37, 38, 39, 36].
As predicted, simulations also show high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.4), consistent with observations that neural responses in primate visual cortex are high-dimensional when visual stimuli are high-dimensional [13].
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)$. As a result, the cancellation mechanism illustrated in Figure 1g manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 3f), a phenomenon that is widely observed in neural recordings [6, 7, 8, 9, 10, 11, 12] and widely studied in computational models [32, 33, 39, 3]. Therefore, the widely studied phenomenon of excitatory-inhibitory balance emerges as a consequence of low-rank suppression.
Amplification of disordered perturbations in networks with spatial structure
Figure 4: Low-rank suppression in a network with spatial structure. a) Connectivity matrix, $W$. Connection strength is a Gaussian function of distance. b) Singular values of $W$ demonstrate effective low-rank structure. c) A spatially smooth perturbation (red) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (gray) is not. d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
Many networks in nature exhibit connection strength that depends on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 34]. As a simple example, we considered a model in which each node is assigned a location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are more strongly connected (Figure 4a). Connectivity was also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank with the first several singular values proportional to the Fourier coefficients of the Gaussian connectivity kernel while the remaining singular values, inherited from $W_{1}$, are much smaller (Figure 4b). The low-rank part is normal, with left- and right singular vectors equal to the Fourier basis vectors (Supplementary Figure S.5), so the network satisfies our conditions for low-rank suppression and high-dimensional responses.
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low frequency spatial Fourier modes. Therefore, due to low-rank suppression, network responses to spatially disordered perturbations are amplified relative to smooth perturbations (Figure 4c-e; compare tick labels on vertical axes). The network also exhibits high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.6), as predicted.
Low-rank suppression and high-dimensional dynamics in an epidemiological network.
Figure 5: Low-rank suppression and high-dimensional dynamics in a real epidemiological network. a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction. b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
We next consider a real epidemiological network, specifically a network of high school social contacts [40], which was used in a recent theoretical study of low-rank networks [1]. In that study, the authors considered quenched mean-field reduction of the susceptible-infected-susceptible epidemiological model,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (5)
where $g>0$, $\gamma>0$ and $\circ$ denotes element-wise multiplication. In this model, each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [40]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a). In [1], it was shown that the dynamics, ${\bf{z}}(t)$, generated by Eq. (5) on this network are effectively low-dimensional. However, Eq. (5) is completely self contained without any internal noise or external perturbations. Noise and perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the high school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network will produce low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
Simulations of these dynamics indeed demonstrate low-rank suppression (Figure 5a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 5b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
Discussion
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. Moreover, in the presence of high-dimensional noise or perturbations, low-rank networks often produce high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional responses to high-dimensional perturbations.
The fact that stable, strongly low-rank networks produce low-rank suppression might seem obvious in hindsight: Whenever $W_{0}$ is normal, stability implies that $W$ has a large, negative eigenvalues, so dynamics are highly suppressive in the direction of the corresponding eigenvectors (as in Figures 1 and 2b,c). However, when $W_{0}$ is highly non-normal $W_{0}$, the network still produces low-rank suppression even when eigenvalues are small in magnitude (Supplementary Figure S.2), so this argument alone does not explain low-rank suppression.
The fact that low-rank networks produce high-dimensional responses to high-dimensional inputs might also seem obvious in hindsight: Inputs drive the network, so high-dimensional input should drive high-dimensional activity regardless the internal network structure. However, our condition on the alignment of singular vectors shows that this explanation is not sufficient. When left- and right-singular vectors are not aligned ($\sigma_{P}\ll 1$), high-dimensional inputs drive low-dimensional responses (Figure 2f,g and Supplementary Figure S.2).
Our results have implications for networks in nature and applications. For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 17, 4, 18, 19, 20, 3, 5, 24]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural dynamics [41, 42, 43]. However, many of these recordings are made during the presentation of low-dimensional stimuli or motor tasks. A growing number of more recent studies have shown that neural activity is high-dimensional during the presentation of high-dimensional stimuli or tasks [13, 43, 44, 45, 46, 47]. Our results are consistent with the hypothesis that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [13]. Our results are also consistent with observations that neural populations are more sensitive to random perturbations than to perturbations that are aligned with low-dimensional network structure [36].
Beyond neuroscience, our results imply that perturbations or inputs to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space. This result can be used to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equation simulations – except for Figure 5 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 5 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
https://github.com/RobertRosenbaum/HighDimLowDimCode
Acknowledgments
We thank Ashok Litwin-Kumar for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
Supplementary Materials
Figure S.1: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau{\bf{z}}^{\prime}=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure S.2: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 1 except $W_{0}=|c|{\bf{u}}{\bf{v}}^{T}$ where ${\bf{u}}$ and ${\bf{v}}$ are independent random unit vectors. Also, in panel c, we used a perturbation of ${\bf{x}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ as given in Eq. (15)) and in panel e-g, we projected onto ${\bf{v}}$. The network exhibits low-rank suppression in response to perturbations parallel to ${\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ (panel c) even though the connectivity matrix, $W$, does not have strongly negative eigenvalues (minimum real part of the eigenvalues of $W$ here is $\textrm{re}(\lambda)_{\textrm{min}}=-0.9$ compared to $\textrm{re}(\lambda)_{\textrm{min}}=-10$ in Figure 1). However, the network produces low-dimensional responses to high-dimensional inputs (panel d).
Figure S.3: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (3), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $-0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure S.4: Low-dimensional dynamics in a modular network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the modular network from Figure 3.
Figure S.5: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 4.
Figure S.6: Low-dimensional dynamics in a spatial network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the spatial network from Figure 4.
Figure S.7: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor ($c=-2$), and the dynamics obey Eq. (3). The instability produced by taking $\rho>1$ generates chaotic dynamics [16, 3].
Figure S.8: Dynamics of an unstable, non-chaotic network with rank-one structure. Same as Figure 1 except $c=10$ and the dynamics obey Eq. (3). The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
Figure S.9: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1d–g except ${\bf{x}}(t)$ was made low-dimensional by multiplying it by a matrix with rank-one structure, ${\bf{x}}\leftarrow W_{x}{\bf{x}}$ where $W_{x}$ is a matrix generated identically to, but independently from $W$. The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$.
Figure S.10: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of ${\bf{z}}$ for the model in Figure 5.
S.1 Comparison to previous work on dynamics of low-rank networks
In this section we review the relationship between our models and analysis and the models and analyses from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 17, 4, 18, 19, 20], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
All three studies quoted above consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model. We next describe the salient properties that distinguish these models from ours.
Thibeault et al. [1] assume that their networks are self-contained and do not receive any time varying external input. Specifically, they explicitly restrict their dynamics to models of the form
$\frac{d{\bf{z}}}{dt}=F({\bf{z}},W{\bf{z}})$
which excludes the possibility of modeling a time-varying external input, like ${\bf{x}}(t)$ in our model. Models of this form cannot describe networks that are part of a larger network, or networks that are modulated by time-varying, external factors. Similar assumptions were made in one study by Landau and Sompolinsky [5]. In other work by Landau and Sompolinsky [3], external input was included, but this input was assumed to be perfectly aligned to the low-rank structure of the connectivity matrix and therefore low-dimensional.
The salient differences between our model and the models considered by Ostojic and colleagues [2, 17, 4, 18, 19, 20] are more subtle. Like us, they consider external inputs that are not aligned to the low-rank part of $W$. Also, like us, they take $W=W_{0}+W_{1}$ where $W_{1}$ is full rank with random entries and the variance of the entries in $W_{1}$ scale like $\mathcal{O}(1/N)$ so that the maximum singular value of $W_{1}$ scales like $\mathcal{O}(1)$.
However, in contrast to our models, the low-rank components of the networks considered by Ostojic and colleagues take the form
$W_{0}=\sum_{\mu=1}^{r}\frac{{\bf{m}}_{\mu}{\bf{n}}_{\mu}^{T}}{N}$ (6)
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (6), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) scale the same in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
This difference in scaling is acknowledged by Ostojic et al. [2, 17, 4, 18, 19, 20] who refer to their networks as “weakly low-rank.” Keeping with this terminology, we will use “weakly low-rank” in this section to refer to networks in which the singular values of the low-rank components ($W_{0}$) and high-rank components ($W_{1}$) both scale like $\mathcal{O}(1)$. We additionally use “strongly low-rank” to refer to networks like ours in which the singular values of $W_{0}$ diverge to $\infty$ as $N\to\infty$ (typically, they scale like $\mathcal{O}(\sqrt{N})$) while the singular values of $W_{1}$ scale like $\mathcal{O}(1)$.
To compare weakly and strongly low-rank networks, we can consider the simple rank-one example from Figure 1. Specifically,
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
while $W_{1}$ has unbiased random entries with variance $\rho/\sqrt{N}$ and dynamics obey Eq. (1). Strongly low-rank networks have $c\gg 1$ while $c\sim\mathcal{O}(1)$ in weakly low-rank networks. In strongly low-rank networks, stability requires that $c<0$, which is critical for low-rank suppression. In weakly low-rank networks, however, it is possible to take $c>0$ without causing instability if $c<1$. When $c>0$, weakly low-rank networks can respond to perturbations aligned to ${\bf{u}}$ more strongly than random perturbations (Supplementary Figure S.11), reversing the trend of low-rank suppression. When $c<0$ in weakly low-rank networks, perturbations in the direction of ${\bf{u}}$ can be suppressed relative to random perturbations, but the effect is weaker than in strongly low-rank networks (Supplementary Figure S.12).
Landau and Sompolinsky [5, 3] consider strongly low-rank networks. However, as noted above, their external input, ${\bf{x}}(t)$, is aligned to the low-rank part of $W$ and is therefore low-dimensional itself. Thibeault et al. [1] consider several network models, but their “rank-perturbed Gaussian” model is strongly low-rank and equivalent to the network structure we study in Figure 1. Complicating matters, Thibeault et al. directly compare their rank-perturbed Gaussian model to the networks in Ostojic et al., despite the fact that the scaling of their low-rank parts differ by a magnitude of $\sqrt{N}$. Specifically, in Thibeault et al. [1], the low-rank part is defined by Eq. (6) where ${\bf{m}}$ and ${\bf{n}}$ are unbiased Gaussian random vectors. The entries of ${\bf{m}}$ have $\mathcal{O}(1/N)$ variance while the entries of ${\bf{n}}$ have $\mathcal{O}(1)$ variance, so the entries of $W_{0}$ have $\mathcal{O}(1/N)$ variance, in contrast to the $\mathcal{O}(1/N^{2})$ variance used by Ostojic et al. [2]. Importantly, this means that the singular values of $W_{0}$ are $\mathcal{O}(\sqrt{N})$ in the rank perturbed Gaussian model analyzed by Thibeault et al. [1], but $\mathcal{O}(1)$ in the models by Ostojic et al. Hence, the rank perturbed Gaussian models considered by Thibeault et al. are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except $c=0.75$.
Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=-0.75$.
S.2 Mathematical analysis of low-rank networks with external perturbations
In this section we present a mathematical analysis of low-rank suppression and high-dimensional dynamics in networks with effective low-rank connectivity and external perturbations. Our analysis applies to a broad class of dynamics. We begin by reviewing the precise assumptions made in our analysis.
S.2.1 Assumptions on the dynamics and definition of the effective connectivity matrix.
We start by assuming dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},{\bf{x}})$ (7)
for some smooth ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$ where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and ${\bf{x}}(t)$ is an $N\times 1$ vector of external forces, inputs, and/or noise. At the end of Section S.2.3, we show how our approach can be generalized to a larger class of causal, time invariant systems, but the dynamics in Eq. (7) are sufficient for all of the examples studied in this manuscript, and the majority of models studied in the literature on low-rank networks, so we focus on this class of dynamics for simplicity.
In general, it is not immediately clear how to define a network of interactions, $W$, from dynamics as general as those in Eq. (7). For example, in nonlinear systems, the effective connectivity between nodes can change with the state of the system. To circumvent this problem, we define the effective connectivity matrix, $W$, through a linearization around a stable equilibrium.
Specifically, we assume that there exists a time-constant forcing, ${\bf{x}}(t)={\bf{x}}_{0}$, that produces a stable steady state solution, ${\bf{z}}_{0}$. In other words,
${\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})={\bf{0}}.$
We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
$J_{{\bf{z}}}=\partial_{{\bf{z}}}F({\bf{z}}_{0},{\bf{x}}_{0})$
has eigenvalues with strictly negative real part. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
${\bf{x}}_{p}(t)={\bf{x}}_{0}+\epsilon{\bf{x}}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, ${\bf{x}}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}+\mathcal{O}(\epsilon^{2}).$ (8)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}(t)$ (9)
where $J_{{\bf{x}}}=\partial_{{\bf{x}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ is the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{x}}$ evaluated at the fixed point, and
$W=J_{{\bf{z}}}+I$ (10)
with $I$ the identity matrix and $J_{{\bf{z}}}=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{z}}$ evaluated at the fixed point. We interpret $W$ in Eq. (10) as the effective connectivity matrix.
It may seem counterintuitive at first to define $W=J_{{\bf{z}}}+I$ instead of $W=J_{{\bf{z}}}$. To understand our choice, note that when we choose $W=J_{{\bf{z}}}+I$, the perturbed system ${\bf{z}}_{p}(t)$ decays to the unperturbed fixed point, ${\bf{z}}_{0}(t)$, whenever there is no perturbation or connectivity (${\bf{x}}(t)=0$ and $W=0$). Moreover, when there is a perturbation without connectivity ($W=0$ and ${\bf{x}}\neq 0$), the perturbation converges to ${\bf{z}}=J_{{\bf{x}}}{\bf{x}}$. Hence, $W$ represents interactions that determine the dynamics under perturbations away from the fixed point. In any case, our overall results are not sensitive to this choice.
One caveat to our approach for defining effective connectivity, $W$, from the dynamics in Eq. (7) is that the input perturbation, ${\bf{x}}(t)$, must be weak enough and/or the dynamics must be sufficiently linear for the linearized dynamics to be accurate, i.e., for the $\mathcal{O}(\epsilon^{2})$ term in Eq. (8) to be small. Moreover, the dynamics of the unperturbed system must be stable near the steady state. Examples with unstable, nonlinear dynamics are considered in Supplementary Figures S.7 and S.8.
Another caveat is that multiplication of ${\bf{x}}$ by $J_{{\bf{x}}}$ in Eq. (9) could suppress dimensionality, for example if $J_{{\bf{x}}}$ is low-rank or effectively low-rank. This would correspond to an effectively feedforward “read-in” mechanism that lowers dimensionality. Specifically, the “effective input” $J_{{\bf{x}}}{\bf{x}}$ would be low-dimensional and suppressed in some directions. For example, consider dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+W_{x}{\bf{x}}$ (11)
where $W_{x}$ is some matrix. In this model, we have $J_{{\bf{x}}}=W_{x}$. Therefore, if $W_{x}$ were low-rank then the effective input, ${\bf{x}}_{eff}=W_{x}{\bf{x}}$, would be low-dimensional even when the raw input, ${\bf{x}}$, were high-dimensional. This scenario would promote low-dimensional dynamics in ${\bf{z}}(t)$. An example of this form is considered in Supplementary Figure S.9. In many models (including Eqs. (1) and (3)), the dynamics take the form ${\bf{F}}({\bf{z}},{\bf{x}})={\bf{F}}({\bf{z}})+{\bf{x}}$ so that $J_{{\bf{x}}}=I$ and this is not an issue. More generally, we assume that $J_{{\bf{x}}}$ is full rank in all of our analysis.
Before continuing, we compare our approach for relating dynamics to connectivity with the approach taken in Thibeault et al. [1]. They define dynamics that explicitly take the connectivity matrix into account,
$\frac{d{\bf{z}}}{dt}$ $={\bf{G}}({\bf{z}},{\bf{y}})$
${\bf{y}}$ $=U{\bf{z}}$
where $U$ is interpreted as the connectivity matrix (they use “$W$”, but we switched to “$U$” to avoid confusion with our definition). This formulation represents a self-contained network in the absence of external perturbations or noise, but it can easily be extended to allow external perturbations by defining
$\frac{d{\bf{z}}}{dt}={\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}}).$ (12)
However, Eq. (12) is not suitable for our purposes because $U$ does not necessarily capture all of the interactions between nodes. For example, the $j$th entry of ${\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}})$ could depend explicitly (through its first argument) on the $k$th entry of ${\bf{z}}$ for some $j\neq k$. This would represent an interaction between nodes $j$ and $k$, which are not accounted for in $U$. For example, we could have $d{\bf{z}}_{j}/dt={{\bf{G}}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{z}}_{k}+{\bf{y}}_{j}={\bf{z}}_{k}+U_{ji}{\bf{z}}_{i}$ for three distinct indices $i,j,k$ where we assume that $U_{jk}=0$. In this case, the connectivity matrix, $U$, correctly captures the influence of node ${\bf{z}}_{i}$ on node ${\bf{z}}_{j}$ through the entry $U_{ji}$. However, node ${\bf{z}}_{j}$ is also influenced by node ${\bf{z}}_{k}$ and this interaction is not captured by the matrix $U$ (as we take $U_{jk}=0$) because it is instead captured by the explicit dependence of ${\bf{G}}$ on ${\bf{z}}_{k}$ (the first argument of ${\bf{G}}$). This could, of course, be overcome by incorporating the dependence on ${\bf{z}}_{k}$ into $U$ (i.e., by setting $U_{jk}=1$), but Eq. (12) makes no assumption that all such interactions are included into $U$. More importantly, it might not be possible to account for higher order, nonlinear interactions using the formulation in Eq. (12). For example, we could have $d{\bf{z}}_{j}/dt={\bf{G}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{y}}_{k}{\bf{z}}_{i}=U_{jk}{\bf{z}}_{k}{\bf{z}}_{i}$ for three distinct indices $i,j,k$. In this case, the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ would not be accounted for in the matrix $U$, and it is not possible to account for both interactions ($i\to j$ and $k\to j$) in the matrix $U$ using the formulation in Eq. (12). Moreover, the strength of the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ depends on the magnitude of ${\bf{z}}_{k}$, which can change with the state of the network. These matters are handled in the mathematical analysis of Thibeault et al. [1] by bounding norms related to the Jacobian matrix of ${\bf{G}}$ with respect to ${\bf{z}}$ and ${\bf{y}}$, and they also separately consider higher order interactions. Our approach avoids these bounds on the Jacobian matrices at the expense of relying on a linearization near a steady state. Under the formulation from Eq. (12), if there is a fixed point at ${\bf{z}}_{0}$ in response to static input ${\bf{x}}_{0}$, then our interpretation of the connectivity matrix is given by
$W=\partial_{{\bf{z}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})+\partial_{{\bf{y}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})U+I$
where ${\bf{y}}_{0}=U{\bf{z}}_{0}$ and the two partial derivatives denote Jacobian matrices with respect to the first and second arguments, respectively. Hence, the two approaches can easily be directly related.
In summary, we consider the dynamics defined by Eq. (7), linearized around a stable steady state, which ultimately gives rise to the definition of network dynamics in Eq. (9). This approach has some caveats, which are discussed above.
S.2.2 Assumptions on the network structure.
The network represents $N$ interacting nodes, the interactions are defined by an $N\times N$ connectivity matrix, $W$, and we consider asymptotics in the $N\to\infty$ limit. Here, we specify the assumptions made about the structure of $W$ in the large $N$ limit.
Letting $\sigma_{1},\ldots,\sigma_{N}$ be the singular values of $W$, we assume that $\sigma_{k}\gg 1$ for $k=1,\ldots,r$ (in the sense that $\sigma_{k}\to\infty$ as $N\to\infty$) and that $\sigma_{k}\leq\mathcal{O}(1)$ for $k=r+1,\ldots,N$ (in the sense that $\sigma_{k}$ has a finite limit as $N\to\infty$). We also assume that $r\sim\mathcal{O}(1)$ so that $r\ll N$. In summary, $W$ has a small number of large singular values and the remaining singular values are much smaller.
Under these assumptions, $W$ can be decomposed as [1]
$W=W_{0}+W_{1}$
where $W_{0}$ has rank $r$ and we write its singular value decomposition as
$W_{0}=U\Sigma V^{T}.$
Here, $U$ and $V$ are $N\times r$ orthonormal matrices and $\Sigma$ is an $r\times r$ diagonal matrix with diagonal entries, $\Sigma_{kk}=\sigma_{k}$. The left and right singular matrices, $U$ and $V$, can be written as
$U=\left[{\bf{u}}_{1}\;\;{\bf{u}}_{2}\;\cdots\;{\bf{u}}_{r}\right]$
and
$V=\left[{\bf{v}}_{1}\;\;{\bf{v}}_{2}\;\cdots\;{\bf{v}}_{r}\right]$
where their columns, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$, are the left and right singular vectors, which satisfy $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$ and ${\bf{u}}_{k}\cdot{\bf{u}}_{j}={\bf{v}}_{k}\cdot{\bf{v}}_{j}=0$ for $j\neq k$. We further assume that $W_{1}$ is a random matrix, which is almost surely full rank and is statistically independent from $W_{0}$. Finally, we assume that the eigenvalues and singular values of $W_{1}$ are smaller than $1$ in magnitude and bounded away from $1$ ($\sigma_{W_{1}},|\lambda_{W_{1}}|<1-\delta$ for some fixed $\delta>0$) with probability $1$.
S.2.3 Analysis of low-rank suppression.
We now analyze the conditions under which low-rank suppression occurs. Our goal here is to understand the conditions under which the dynamics produce low-rank suppression like that observed in Figure 1c. In the next section, we will derive conditions on high-dimensional dynamics in the presence of stochastic perturbations such as those considered in Figure 1d–g.
We begin by considering the case in which ${\bf{x}}(t)$ is static or slow, so we can use a steady state or quasi-steady state approximation. Later in this section, we consider low-rank suppression with transient and fast perturbations.
If ${\bf{x}}(t)={\bf{x}}$ is static (such as during the duration of each stimulus in Figure 1c), then the network response converges to solutions of the perturbed steady state, which from Eq. (9) satisfies
${\bf{z}}=W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}$
and therefore
${\bf{z}}=[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}.$ (13)
If ${\bf{x}}(t)$ is not static, but changes much more slowly than $\tau$, then ${\bf{z}}(t)$ closely tracks the quasi-steady state given by Eq. (13).
Low-rank suppression, like that observed in Figure 1c, occurs whenever there is an input, ${\bf{x}}$, that is strongly suppressed by the recurrent dynamics in comparison to a random input (as in Figure 1c). In other words, we can say that low-rank suppression occurs whenever there is a raw input, ${\bf{x}}$, such that $\|{\bf{x}}\|=\mathcal{O}(1)$ and
$\|{\bf{z}}\|=\|[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}\|\ll\|{\bf{z}}_{rand}\|=\|[I-W]^{-1}{\bf{u}}_{rand}\|$
with high probability whenever ${\bf{u}}_{rand}$ is a random unit vector, independent from $W_{0}$, $W_{1}$, and $J_{{\bf{x}}}$. To simplify this statement, first note that $\|[I-W]^{-1}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ whenever ${\bf{u}}_{rand}$ is independent from $W_{1}$. Moreover, since we assume that $J_{{\bf{x}}}$ is full rank, we can write the condition in terms of the effective input,
${\bf{x}}_{eff}=J_{{\bf{x}}}{\bf{x}}$
instead of the raw input, ${\bf{x}}$.
In particular, low-rank suppression occurs whenever there is a vector ${\bf{x}}_{eff}$ such that
$\|{\bf{z}}\|=\|[I-W]^{-1}{\bf{x}}_{eff}\|\ll\|{\bf{x}}_{eff}\|.$ (14)
We take this property to be the definition of low-rank suppression.
For the dynamics that we use in our simulations, note that $J_{{\bf{x}}}=I$ so ${\bf{x}}_{eff}={\bf{x}}$. We discuss some caveats of the distinction between ${\bf{x}}$ and ${\bf{x}}_{eff}$ at the end of this section. We now prove that low-rank suppression arises under the assumptions from Sections S.2.2S.2.3.
Claim 1.
Under the model and assumptions of Sections S.2.2S.2.3, the network exhibits low-rank suppression.
Proof.
Let $k\in\{1,\ldots,r\}$ and take
${\bf{x}}_{eff}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (15)
We claim that the response to this input is given by
${\bf{z}}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
To see this, we only need to show that $[I-W]{\bf{z}}={\bf{x}}$. We have
${\bf{z}}$ $=\frac{W_{0}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$=\frac{U\Sigma V^{T}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{x}}_{eff}.$
Now note that $\|{\bf{x}}_{eff}\|=1+o(1)$ and $\|{\bf{z}}\|\ll 1$ because $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, $\|W_{1}\|_{2}=\mathcal{O}(1)$, and $\sigma_{k}\gg 1$. Therefore $\|{\bf{z}}\|\ll\|{\bf{x}}_{eff}\|$. ∎
Before continuing, we will make a few observations about this claim and its proof. First note that the definition of low-rank suppression in Eq. (14) is equivalent to the statement that the matrix
$A=[I-W]^{-1}$
has at least one asymptotically small singular value,
$\sigma_{A}\ll 1.$
Singular values commute with matrix inverses, so the singular values of $A$ are given by
$\sigma_{A}=\frac{1}{\sigma_{Q}}$
where $\sigma_{Q}$ are the singular values of the matrix
$Q=I-W.$
Therefore, low-rank suppression occurs whenever $Q$ has at least one asymptotically large singular value,
$\sigma_{Q}\gg 1.$
Since $W$ has an asymptotically large singular value, low-rank suppression occurs because this large singular value is conserved by adding the identity matrix. However, unlike eigenvalues, singular values are not transformed in a simple way by adding the identity matrix, in other words, we do not generally have $\sigma_{Q}=1-\sigma_{W}$. If $W$ were a normal matrix, then we would have $\sigma_{Q}=|1-\sigma_{W}|$, so the emergence of low-rank suppression would have been easier to prove, but this is not in general true of non-normal matrices. The potentially complicated relationship between the singular values of $W$ and those of $Q$ is more critical for the analysis of high-dimensional dynamics in the next section.
The proof of Claim 1 also provides insight into the input and response directions that are suppressed in low-rank suppression. Note that ${\bf{x}}_{eff}={\bf{u}}_{k}+o(1)$ and ${\bf{z}}\propto{\bf{v}}_{k}$, which implies that inputs approximately aligned to the column space of $U$ (equivalently, the column space of $W_{0}$) give rise to suppressed responses aligned to the column space of $V$ (equivalently, the row space of $W_{0}$).
Finally, it is worth discussing our use of ${\bf{x}}_{eff}$ instead of ${\bf{x}}$ in the definition of low-rank suppression. For our simulations ${\bf{x}}={\bf{x}}_{eff}$ so there is no issue. However, one could consider a model for which $J_{{\bf{x}}}$ is full rank, but amplifies some directions more than others. For example, consider the model in Eq. (11) with a $W_{x}$ chosen to amplify inputs in every direction parallel to the column space of $W$, i.e.., if $\|W_{x}{\bf{u}}_{k}\|\gg 1$ for $k=1,\ldots,r$ whereas $\|W_{x}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ for random directions. In this very specific case, the amplification by $W_{x}$ would compete with the suppression by $A=[I-W]^{-1}$, so the raw input, ${\bf{x}}$, would need to be smaller in the direction of ${\bf{u}}_{k}$ to obtain an ${\bf{x}}_{eff}$ that is $\mathcal{O}(1)$ in that direction. In this case, the suppression by recurrent dynamics would compete with the amplification by a feedforward read-in. Importantly, the recurrent dynamics would still be suppressive to effective inputs aligned to ${\bf{u}}_{k}$. We focus on effective inputs in our definition of low-rank suppression to avoid such an issue, but it would only arise in a very specific situation where $J_{{\bf{x}}}$ amplifies exactly the directions that $A=[I-W]^{-1}$ suppresses.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-static approximation from Eq. (13) in this case. Instead, we take the Laplace transform in Eq. (9) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (16)
where
$\hat{H}(s)$ $=(1+\tau s)^{-1}\left[J_{{\bf{z}}}+I\right]=(1+\tau s)^{-1}W$ (17)
$\hat{B}(s)$ $=(1+\tau s)^{-1}J_{{\bf{x}}}$
[…]
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$, i.e.,
$W=\hat{H}(s)$
because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. Solving the implicit Eq. (16) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-W]^{-1}\hat{B}\hat{{\bf{x}}}.$ (18)
Since $\hat{B}$ quantifies feedforward, read-in dynamics, we can interpret $\hat{B}\hat{{\bf{x}}}$ as effective input in the Laplace domain, ${\bf{\hat{x}}}_{eff}=\hat{B}\hat{{\bf{x}}}$. This is analogous to our treatment of $J_{{\bf{x}}}{\bf{x}}$ above.
Eq. (18) is identical to Eq. (13) except that Eq. (18) quantifies interactions at individual Laplace modes, $s$. Hence, Claim 1 applies to situations in which ${\bf{x}}(t)$ is transient and fast, so long as the effective connectivity, $W$, and effective input, ${\bf{x}}_{eff}$ are interpreted in the Laplace domain ($W=H(s)$ and ${\bf{x}}_{eff}=\hat{B}(s)\hat{{\bf{x}}}(s)$).
Moreover, the approach of working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (7). The dynamics in Eq. (7) capture a large class of models, but are not fully general. Specifically, in Eq. (7), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$ and the values of $x(s)$ for $s\leq t_{1}$. However, some systems like integro-differential equations and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (7). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (19)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivriant function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (19) encompasses an extremely broad class of systems, including those described by Eq. (7).
We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq. (19). As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$. However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$. We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
A first order Volterra expansion gives Eq. (16) as an implicit equation for ${\bf{z}}(t)$. The matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (9), we obtain the $\hat{H}$ and $\hat{B}$ defined in Eq. (17). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (19).
S.2.4 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–g. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
[…]
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. This is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (9) to get [48, 49, 34, 35, 50]
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}J_{{\bf{x}}}\widetilde{C}^{{\bf{x}}}(f)J_{{\bf{x}}}^{T}[\widetilde{r}(f)I-W]^{-*}$ (20)
where
$\widetilde{r}(f)=1-2\pi\tau fi$
is scalar.
Note that Eq. (20) can be derived from Eqs. (18) and (17) by first taking $s=-2\pi if$ in Eq. (18) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the results derived in this section apply to the general class of dynamics defined in Eq. (19).
Eq. (20) quantifies the covariance structure of ${\bf{z}}(t)$ at any given frequency mode, $f$, but we are often specifically interested in the zero-lag temporal covariance,
[…]
From Eq. (20), we therefore have
[…]
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (21) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}J_{{\bf{x}}}\overline{C}^{{\bf{x}}}J_{{\bf{x}}}^{T}[I-W]^{-T}.$ (22)
When principal component analysis is applied to ${\bf{z}}(t)$, the variance explained by each principal component is given by the ordered list of eigenvalues of the covariance matrix, $\overline{C}^{{\bf{z}}}$. From Eq. (22), we see that these eigenvalues are proportional to the squares of the singular values of the matrix
$R=[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}.$
where $\sqrt{\overline{C}^{{\bf{x}}}}$ is the matrix square root of the symmetric positive definite covariance matrix, $\overline{C}^{{\bf{x}}}$ (distinct from the entry-wise square root in general). Therefore, the decay of the variance explained by each principal component of ${\bf{z}}$ (as in the blue dots in Figure 1d) are described by the squared singular values of $R$.
In the models we consider, each ${\bf{x}}_{j}(t)$ is an i.i.d. stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$. Moreover, in the networks we consider, $J_{{\bf{x}}}=I$ is equal to the identity matrix. Therefore, for our models,
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}v$
where $v$ is a scalar.
In more general classes of models in which $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are not multiples of the identity matrix, dimensionality of the dynamics of ${\bf{z}}(t)$ could be reduced whenever $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) or a low-rank read-in mechanism ($J_{{\bf{x}}}$ low-dimensional; see Supplementary Figure S.9 and discussion in previous sections), and would not be related to the actual dynamics of interactions within the network. Therefore, our focus on the effective rank of $[I-W]^{-1}$ in place of $[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}$ is justified even in more general settings.
In conclusion, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
$A=[I-W]^{-1}.$
More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
[…]
where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
Therefore, ${\bf{z}}(t)$ is low-dimensional whenever $A$ has a small number of large singular values. As above, we can use the fact that singular values commute with matrix inverses to write our conclusions in terms of
$Q=I-W.$
Specifically,
[…]
where $\sigma_{Q,N-k}$ is the $(N-k)$th singular value of $Q$, i.e., the $k$th from the last singular value (assuming singular values are sorted in decreasing order).
Figure S.13: Comparison between theory and simulations for shorter and longer simulations. a) Same as Figure 1d except we added the theoretical values form Eq. (23) (red dots) and the values obtained from Eq. (24) (green dots). The network was simulated for $T=5\times 10^{3}$ time units. b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (23) as red dots (Supplementary Figure S.13a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (22) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (23) would be replaced by
[…]
where
$U=[I-W]\sqrt{{\overline{C}}^{\bf{x}}}^{-1}$
and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
[…]
where
$\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
is the sampled value of $R$ (assuming $J^{{\bf{x}}}=I$) and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (24) gives a much more accurate prediction (Supplementary Figure S.13a, green dots) This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (23) was accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
As concluded at the end of the previous section, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those reached in Section S.2.3, we can summarize as follows
Low-rank suppression (as in Figure 1c) occurs whenever $Q=I-W$ has at least one asymptotically large singular value.
High-dimensional dynamics (as in Figure 1d) occur whenever $Q=I-W$ lacks any asymptotically small singular values.
In the previous section, we showed that $Q=I-W$ has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which $Q=I-W$ lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
$P=V^{T}U,$
which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$. Specifically, $P_{jk}={\bf{v}}_{k}\cdot{\bf{u}}_{k}$ so that $P_{jk}=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$. Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$. We next show that if all singular values of $P$ are $\mathcal{O}(1)$ then the network exhibits high-dimensional responses to high-dimensional inputs.
Claim 2.
Under the model and assumptions of Sections S.2.2 and S.2.4, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
Proof.
We will prove this claim by proving its contrapositive:
If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever $Q=I-W$ has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:
If $Q=I-W$ has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
We will prove this version of the claim directly. Assume that $Q$ has an asymptotically small singular value. Then there is a ${\bf{z}}$ satisfying $\|{\bf{z}}\|=1$ and
$(I-W){\bf{z}}=o(1)$
where the notation $o(1)$ means that $\|(I-W){\bf{z}}\|\to 0$ as $N\to\infty$. It is sufficient to show that there is a ${\bf{y}}$ with $\|{\bf{y}}\|=1+o(1)$ satisfying
$\|P{\bf{y}}\|=o(1).$
We have that
$W_{0}{\bf{z}}+W_{1}{\bf{z}}={\bf{z}}+o(1)$
Multiplying both sides on the left by $U^{T}$ gives
$\Sigma V^{T}{\bf{z}}+U^{T}W_{1}{\bf{z}}=U^{T}{\bf{z}}+o(1)$
Since $W_{1}$ is random and independent from $W_{0}$ and $U$, and since $U^{T}$ projects from $N$ to $r\ll N$ dimensions, the term $U^{T}W_{1}{\bf{z}}$ represents the projection of a random vector onto the low-dimensional column space of $U$, so $\|U^{T}W_{1}{\bf{z}}\|=o(1)$. We therefore have
$\Sigma V^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
We next claim that ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. To see why this is true, note that $UU^{T}{\bf{z}}$ is the orthogonal projection of ${\bf{z}}$ onto the column space of $W_{0}$. Since $W_{0}$ dominates $W=W_{0}+W_{1}$ (and therefore the the column space of $W$ is dominated by that of $W_{0}$) and $W{\bf{z}}\approx{\bf{z}}$, we may conclude that ${\bf{z}}$ lies predominantly in the column space of $W$ and therefore of $W_{0}$. In other words, ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. Hence, we can rewrite the equality above as
$\Sigma V^{T}UU^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
which reduces to
$\Sigma P{\bf{y}}={\bf{y}}+o(1).$
where $P=V^{T}U$ and ${\bf{y}}=U^{T}{\bf{z}}$. Note again that $\|{\bf{y}}\|=\|U^{T}{\bf{z}}\|=1+o(1)$ since $\|{\bf{z}}\|=1$ and ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. We then have that
$\|P{\bf{y}}\|=\|\Sigma^{-1}{\bf{y}}\|+o(1)=o(1)$
because $\Sigma^{-1}$ is a diagonal matrix with $o(1)$ terms on the diagonal. ∎
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2g. Since ${\bf{z}}$ is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$. Indeed, in Figure 2g, the angle between ${\bf{u}}_{2}$ and the dominant principle component direction of ${\bf{z}}$ is $18^{\circ}$.
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (20) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (22). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (20) instead of the identity itself, as in Eq. (22). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).