### v4

# High-dimensional dynamics in low-dimensional networks

Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive novel, precise conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
Introduction
Recent work shows that many networks arising in nature and applications have low-dimensional structure [1]. This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15]. However, many of these recordings are made in the context of low-dimensional stimuli or behavior. A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20]. These observations raise the important point that the dimensionality of dynamics on a network likely depends on the structure of the network’s external input, in addition to the network’s internal structure.
Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [1, 5, 12]. Other work considers external input, but assumes that the input is perfectly aligned with the network’s low-dimensional structure [3]. In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure. Other work [2, 4] considers the case of general external input, but the recurrent network is assumed to be weakly low-rank in the sense that the largest singular values are not asymptotically larger than the bulk of the singular values.
Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk. We show that such networks can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the detailed structure of the network connectivity. Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.”
We use mathematical analysis to derive precise conditions under which networks with low-rank structure produce high-dimensional dynamics and low-rank suppression. Our analysis shows that low-rank suppression is a general feature of networks with linearly stable dynamics and “strong” low-rank structure in the sense that the dominant singular values of the network are large.
On the other hand, high-dimensional dynamics are less general, and only arise when the network’s connectivity matrix has a non-vanishing component that satisfies the equal projector (EP) property. To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been pointed out in any previous work.
We show that many common structural features such as biased weights, modularity, and spatial connectivity naturally produce connectivity matrices with the EP property, implying that these networks produce high-dimensional dynamics in addition to low-rank suppression. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
Our conclusions have important implications for the interpretation of low-dimensional network structure. In neuroscience, our results can explain why neural populations generate high-dimensional responses to high-dimensional stimuli and tasks [16]. Our results also generalize and extend the theory of excitatory-inhibitory balance [21, 3], and amplification arising from breaks in this balance [22, 23, 24]. Beyond neuroscience, our results imply that perturbations misaligned to low-rank network structure are most effective at driving responses. This counterintuitive observation could be used to design more effective interventions to epidemiological, biological, social, and other networks.
Results
High-dimensional dynamics and low-rank suppression in a network with rank-one structure
For illustrative purposes, we begin with a simple, linear model (Figure 1a)
$\tau\frac{d{\bf{z}}}{dt}$ $=-{\bf{z}}+W{\bf{z}}+{\bf{x}}.$ (1)
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
$W=W_{0}+W_{1}$
where
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
is a rank-one matrix and
$W_{1}=\frac{\rho}{\sqrt{N}}Z$
is a full rank random matrix. Specifically, $\bf{u}$} is a random vector with $\|{\bf{u}}\|=1$, $Z$ is an $N\times N$ matrix with entries drawn i.i.d. from a standard normal distribution, and $\rho>0$. If
$|c|\gg\rho$
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (bounded by $2\rho$; Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [25, 26].
Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [27]. If the dynamics were unstable, then $\|{\bf{z}}(t)\|$ would grow exponentially toward $\infty$, so we only consider parameter regimes with stable dynamics. Since strong low-rank structure also requires $|c|\gg\rho$, stability requires $c<0$ when $\rho=\mathcal{O}(1)$. In simulations here, we take $\rho=0.5$ and c=10c=-10. We consider more general network structures later.
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 10 in Appendix D). However, this conclusion does not necessarily carry over to recurrent networks.
In the recurrent network, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$. The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d). Perhaps surprisingly, this weakest principal component direction was closely aligned to $\bf{u}$} (the angle was less than $8^{\circ}$). Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of $\bf{u}$} was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The eigenvalues of $W$ have a dominant, negative term. d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}. g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
To better understand these results, we next simulated the network with two different external input patterns. One pattern was aligned with the low-dimensional structure of $W$,
${\bf{x}}_{\textrm{aligned}}={\bf{u}},$
while the other had a random direction,
${\bf{x}}_{\textrm{random}}={\bf{u}}_{\textrm{rand}},$
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 10 in in Appendix D). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
We use the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to the low-rank structure of a network are suppressed by the network’s dynamics. In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional inputs (as in Figure 1d).
The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In Appendix A, we provide a detailed review of this previous work and its relation to ours.
A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady-state-solution to Eq. (1),
${\bf{z}}=[I-W]^{-1}{\bf{x}}.$ (2)
Convergence to the steady-state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near $\bf{u}$}. Hence, the recurrent network suppresses inputs in the direction of $\bf{u}$}. A related explanation is that the Jacobian matrix J=WIJ=W-I of Eq. (1) has a large, negative eigenvalue near c1<0c-1<0 with associated eigenvector near $\bf{u}$} so the dynamics in Eq. (1) are highly compressive along $\bf{u}$}. This second explanation relies on the assumption that $W_{0}$ is a symmetric matrix (or at least a normal matrix) but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response. In the qausi-steady state,
${\bf{z}}\approx W{\bf{z}}+{\bf{x}}.$ (3)
In other words, $\bf{z}$} tracks the sum of internal and external inputs in the quasi-steady state.
Since $W$ has a large singular value with left- and right singular values aligned to $\bf{u}$}, multiplication by $W$ amplifies the direction $\bf{u}$}. Therefore, the internal input, $W{\bf{z}}$, is large in the direction of $\bf{u}$} whenever $\bf{z}$} is moderate in the direction of $\bf{u}$}. In other words, $|{\bf{u}}\cdot W{\bf{z}}|\gg|{\bf{u}}\cdot{\bf{z}}|$. This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq. (3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}.
This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}. Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h). We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [21, 28] as we will show later.
Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue. It is widely known that an eigenvalue with large, negative real part produces a suppressive dynamic. Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input. We next address these counterpoints with an example in which $W_{0}$ is highly non-normal. In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
Figure 2: Response properties of a non-normal recurrent network with rank-one structure. Same as Figure 1 except the low-rank part of $W$ is non-normal. The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c). The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
We specifically consider an example in which the model and all parameters are the same as in Figure 1 except that
$W_{0}=c{\bf{u}}{\bf{v}}^{T}$
where $\bf{u}$} and $\bf{v}$} are random, orthogonal vectors with $\|{\bf{u}}\|=\|{\bf{v}}\|=1$. This matrix is highly non-normal in the sense that its left- and right-singular vectors are orthogonal, and its non-zero eigenvalue does not have a simple relationship to its singular values.
The distribution of singular values of $W=W_{0}+W_{1}$ are identical to those from our previous example (Figure 2b; compare to Figure 1b), but the eigenvalues are different (Figure 2c; compare to Figure 1c). Specifically, the eigenvalues no longer have a dominant negative outlier. They are all $\mathcal{O}(1)$ in magnitude and clustered in a single bulk.
Repeating the simulations from above for this network, we found that the response to high-dimensional input was low-dimensional in the sense that variability was much greater in one direction than all others: Over 40% of the variance in ${\bf{z}}(t)$ was captured by a single principle component (first blue dot in Figure 2d), compared to less than 2% for any orthogonal direction (the remaining blue dots). This dominant direction was partially aligned to $\bf{u}$} (the angle between the dominant direction and $\bf{u}$} was $31^{\circ}$). Consistent with this observation, variability was $100$ times larger in the direction of $\bf{u}$} than in a random direction (Figure 2e). Hence, low-dimensional dynamics emerged even in the presence of high-dimensional input.
It might be tempting to conclude that non-normal structure implies low-dimensional dynamics, but in the next section we will show that EP matrices, not normal matrices, are the correct class of matrices to constrain the dimensionality of network dynamics.
In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example. This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$). While the variability along $\bf{v}$} itself was not much smaller than a random directions (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f). Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g). The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in a later section, but note that it is closely aligned to $\bf{u}$} whenever $c$ is large. In other words, inputs in the direction of ${\bf{x}}_{\textrm{aligned}}$, which is close to $\bf{u}$}, are suppressed by the network. Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. In the next section, we derive precise conditions for low-rank suppression and low-dimensional dynamics.
General conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple network with rank-one structure. Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and derive precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
Assumption 1: Dynamics are linear or near a stable equilibrium.
We consider a general class of dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},\bm{\eta})$ (4)
for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and $\bm{\eta}(t)$ is an $N\times 1$ vector of external perturbations, inputs, and/or noise.
In general, the dependence of ${\bf{F}}({\bf{z}},\bm{\eta})$ on $\bf{z}$} defines interactions between nodes, but it is not immediately clear how to define a fixed network of interactions, $W$, from the dynamics in Eq. (4). When dynamics are linear, $\bf{F}$} can be written as ${\bf{F}}({\bf{z}},\bm{\eta})=W{\bf{z}}+{\bf{G}}(\bm{\eta})$, and we can take $W$ to represent the effective connectivity. In nonlinear systems, the effective connectivity between nodes can change with the state of the system. To resolve this difficulty in defining network structure when dynamics are nonlinear, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
Specifically, we assume that there exists a time-constant forcing, $\bm{\eta}(t)=\bm{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$. In other words, there exist $\bm{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
${\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})={\bf{z}}_{0}.$
We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$
has eigenvalues with strictly negative real part. Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
$\bm{\eta}_{p}(t)=\bm{\eta}_{0}+\epsilon\bm{\eta}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\bm{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$ (5)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (6)
where
$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ (7)
is interpreted as the effective connectivity matrix and
${\bf{x}}(t)=\partial_{\bm{\eta}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})\bm{\eta}(t)$
is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$. In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq. (4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq. (6), so care must be taken in interpreting our results. For illustrative purposes, we consider a contrived example of such a model in Figure 11 in Appendix D.
Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix. For example, the system might be written as $d{\bf{z}}/dt={\bf{H}}(W{\bf{z}})+{\bf{x}}$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{H}}(W{\bf{z}}_{0})\neq W$. In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality. We consider one such example from epidemiology at the end of the Results.
The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive. However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Figures 13 and 14 in Appendix D for two examples of nonlinear, unstable dynamics). Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state. Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input). This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$. We discuss the relationships between their analysis and ours in more depth in Appendix A.
If a model is fully linear, i.e., if a model is already written in the form of Eq. (6), then we do not need to assume that dynamics remain close to the equilibrium, so inputs and perturbations need not be weak when dynamics are linear.
In summary, we assume that the network dynamics are either fully linear or, if they are nonlinear, they represent a small perturbation away from a stable equilibrium. This assumption allows us to focus our analysis on the linearized dynamics in Eq. (6).
Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
We now formalize an asymptotic notion of a strongly low-rank network. Specifically, we consider the limit of large network size, $N\to\infty$, and assume that $W$ has $r\sim\mathcal{O}(1)$ asymptotically large singular values,
$\sigma_{k}\gg 1,\;\;k=1,\ldots,r$
where $\sigma_{k}$ denotes the $k$th largest singular value of $W$. We assume that the remaining singular values are small or moderate in magnitude,
$\sigma_{j}\leq\mathcal{O}(1),\;\;j=r+1,\ldots,N.$
Later, we consider some examples in which these scaling laws arise naturally.
Matrices that satisfy the assumptions above can be decomposed as [1]
$W=W_{0}+W_{1}$
where $W_{0}$ has rank $r$ and asymptotically large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank. Since $W_{0}$ has rank $r$, we can write its singular value decomposition as
$W_{0}=U\Sigma V^{T}$
where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values ($\Sigma_{kk}=\sigma_{k}\gg 1$) while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
We additionally assume that $W_{1}$ is a random matrix that is independent of $U$ and $V$. In other words, $W$ is a low-rank matrix perturbed by noise. This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
Strongly low rank structure is a special case of low-rank structure. For example, a network could have a small number of $\mathcal{O}(1)$ singular values with the remaining singular values $o(1)$. Such a network would still be effectively low-rank in an asymptotic sense, but it would not be strongly low-rank according to our definition because the dominant singular values are not asymptotically large. These “weakly low-rank” networks [2] can exhibit a weaker form of low-rank suppression and high-dimensional dynamics, depending on the details of their structure (see Appendix A for further discussion and examples).
Low-rank suppression occurs under Assumptions 1 and 2.
Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically,
Claim 1.
Consider the dynamics defined by Eq. (6) under Assumptions 1 and 2. There exists an input ${\bf{x}}_{a}$ with $\|{\bf{x}}_{a}\|=\mathcal{O}(1)$ such that the network’s steady-state response to ${\bf{x}}_{a}$ is asymptotically larger than the $\mathcal{O}(1)$ response to a random input, ${\bf{x}}_{\textrm{r}}$ (where $\|{\bf{x}}_{\textrm{r}}\|=\mathcal{O}(1)$).
To show why Claim 1 is true, first choose a $k\in\{1,\ldots,r\}$ and define
${\bf{x}}_{a}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (8)
Note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$. Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
${\bf{z}}_{a}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (8). In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is normal. This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
Note from Eq. (8) that ${\bf{x}}_{a}$ is closely aligned to ${\bf{u}}_{k}$ since $\sigma_{k}\gg 1$. Hence, the suppressed input, ${\bf{x}}_{a}$, is nearly (but not exactly) parallel to the left singular vector, ${\bf{u}}_{k}$. Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$. Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 10 in Appendix D for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
In summary, recurrent networks with strongly low-rank structure satisfying Assumptions 1 and 2 admit a small number of input directions along which inputs are strongly suppressed. We call this form of suppression low-rank suppression. In later sections, we will explore consequences of low-rank suppression in common low-rank network structures. Before discussing the dimensionality of dynamics, we first need to make an additional assumption about the dimensionality of the input.
Assumption 3: External inputs are high-dimensional and slowly varying.
To begin discussing the dimensionality of ${\bf{z}}(t)$, we need a model of ${\bf{x}}(t)$ that allows us to quantify dimensionality in a concrete way. To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process in which case ${\bf{z}}(t)$, as defined in Eq. (6), is also a stationary ergodic process whenever dynamics are stable [31]. This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Appendix B for more mathematical details).
The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics. When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 11 in Appendix D for an example).
We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions. To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$. We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
In Appendix B, we give a more mathematically detailed definition of this assumption. In Appendix C, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d. Gaussian process with correlation time $\tau_{x}=10$. Our assumption of high-dimensionality is satisfied because each ${\bf{x}}_{j}(t)$ is independent. Our assumption that ${\bf{x}}(t)$ varies slowly is satisfied because $\tau_{x}$ is much larger than the network’s intrinsic timescale, $\tau=1$.
Conditions for low-dimensional dynamics.
In Appendix B, we prove that (under Assumptions 1–3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
$P=V^{T}U$
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [11]. Our main result from Appendix B can be stated as follows:
Claim 2.
Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
Here, “low-dimensional dynamics” is interpreted to mean that the covariance matrix of ${\bf{z}}(t)$ has an $\mathcal{O}(1)$ number of dominant eigenvalues, equivalently, ${\bf{z}}(t)$ has a small number of dominant principal component directions. Note that this result is equivalent to the following:
Claim 2.
Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If all singular values of $P=V^{T}U$ are $\mathcal{O}(1)$ then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal and symmetric. c) The variance explained by each principal component of the network dynamics, ${\bf{z}}(t)$. d) The eigenvalues of $W$. e–g) Same as b–d except $W_{0}$ is EP, but non-normal. h–j) Same as b–d except $W_{0}$ is not EP. k–m) Same as h–j but using a different non-EP structure. n–p) Same as b–d except $W_{0}$ is non-EP with a non-vanishing EP component. See the text for the precise definition of $W_{0}$ in each case.
To better understand the conditions on $P$ in Claim 2, we consider simulations of a widely used model from computational neuroscience [27, 32, 2, 33],
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (9)
We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 3a) and we fix
$U=\left[\begin{array}[]{cc}{\bf{u}}_{1}&{\bf{u}}_{2}\end{array}\right]$
while exploring different choices of $V$.
We first consider a case in which $W_{0}$ is normal and symmetric by setting
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&-{\bf{u}}_{2}\end{array}\right].$
This is an extension of the example from Figure 1 to rank $r=2$. In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c). Like the example from Figure 1, $W_{0}$ has eigenvalues with large, negative real part (Figure 3d). In this example, $W_{0}$ is both normal and symmetric. Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
High-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied.
An example of a non-normal, EP matrix, $W_{0}$, is given by taking
$V$ $=\left[\begin{array}[]{cc}{\bf{u}}_{2}&-{\bf{u}}_{1}\end{array}\right].$
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$. For this example, $P$ is zero along the diagonal (Figure 3e), but all singular values of $P$ are $\sigma_{P}=1$, so the network still produces high-dimensional dynamics (Figure 3f). Note that $W$ does not have any eigenvalues with large, negative real part (Figure 3g). Hence, this example combined with the previous example (Figure 3b–d) show that high-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
To obtain an example of a network that produces low-dimensional dynamics (see also the example in Figure 2), we set
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&{\bf{v}}_{\perp}\end{array}\right].$
where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=1$ and another at $\sigma_{P}=0$ (Figure 3h). Because of the singular value at zero, the condition in Claim 2 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3i). For this example, $W$ has an eigenvalue with large, negative real part (Figure 3j). However, low-dimensional dynamics also arise in the case (Figure 3k,l)
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&{\bf{v}}_{\perp}\end{array}\right].$
in which case $W$ lacks eigenvalues with large, negative real part (Figure 3m). Hence, these two examples demonstrate that low-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
So far, in all of the examples with high-dimensional dynamics, $W_{0}$ is an EP matrix. However, $W_{0}$ need not be fully EP for dynamics to be high-dimensional. We now consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
where $0<c<1$. This is a linear combination of the EP and non-EP matrices from the examples above. In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3n). As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3o). Also, $W$ has eigenvalues with large, negative real part (Figure 3p). From this example, we may conclude that the low-rank part of the connectivity matrix only needs to have a non-vanishing EP component to produce high-dimensional dynamics.
The numerical results in Figure 3 provide an interpretation of Claim 2: If $W_{0}$ has a non-vanishing component with the EP property, then dynamics are high-dimensional. To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been made before.
The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks with high-dimensional input. As indicated above by Claim 1, the eigenvalue spectrum is also not useful for determining whether strongly low-rank networks produce low-rank suppression.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 12 in Appendix D; compare to Figure 2). However, many networks arising in nature do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
Biased weights and modular networks
So far, we considered networks with unbiased weights, $E[W_{jk}]=0$, which is common in modeling studies, but many networks in nature have weights with non-zero mean. Biased weights can produce low-rank structure. As a simple example, consider a random network with independent weights satisfying
$E[W_{jk}]=m\;\textrm{ and }\;\textrm{std}(W_{jk})=s.$
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix D; see Appendix B in [35] for related results).
Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network. a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure. c) The alignment matrix shows that the network is EP, but not normal. d) An input that is constant within each block (red) is aligned to the low-rank part, but a random input (gray) is not. e) Response magnitude is suppressed for the aligned input relative to the random input. f) The excitatory (positive; red) component of an input balances with the inhibitory (negative; blue) component to produce a much smaller total (gray) component, a widely observed phenomenon in neural circuits.
More generally, modular structure arises in many natural settings [1, 36]. Specifically, many networks in nature represent the interaction between $n$ populations, and mean connection weights between populations are often non-zero. The adjacency matrices of modular networks can be arranged to have a block structure
$W=\left[\begin{array}[]{ccc}W^{1,1}&\cdots&W^{1,n}\\ W^{2,1}&\cdots&W^{2,n}\\ \vdots&\cdots&\vdots\\ W^{n,1}&\cdots&W^{n,n}\end{array}\right]$
where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
$E\left[W^{a,b}_{jk}\right]=m_{ab}\;\textrm{ and }\;\textrm{std}\left(W^{a,b}_{jk}\right)=s_{ab}.$
If we assume that each population has $N_{a}\sim\mathcal{O}(N)$ members and that the $m_{ab}$ scale similarly to the $s_{ab}$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(s_{ab}\sqrt{N})$. Hence, when $s_{ab}\sim m_{ab}$, large modular networks with biased weights naturally produce a low-rank structure in the sense that a small number of singular values are asymptotically larger than the rest [1]. If additionally $m_{ab}\gg\mathcal{O}(1/N)$, then the dominant singular values are asymptotically large, so the network is strongly low-rank in the sense defined in Assumption 2.
Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is constant within each block, so $W_{0}$ has rank at most $n$. In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the indicator vectors of the $n$ populations,
$\textrm{col}(U)=\textrm{col}(V)=\textrm{span}\left\{{\bf{1}}_{1},{\bf{1}}_{2},\ldots,{\bf{1}}_{n}\right\}$
where each ${\bf{1}}_{k}$ is an $N$-dimensional indicator vector with entries defined by
${\bf{1}}_{k}]_{j}=\begin{cases}1&\textrm{if index $j$ is in population $k$}\\ 0&\textrm{otherwise}\end{cases}$
Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics.
[…]
Together, these biologically justified assumptions imply that the network has a strongly low rank structure, as defined in Assumption 2, with rank $r=2$ (Figure 4b). Specifically, Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$. Moreover, the low-rank part of the connectivity matrix is EP (Figure 4c). Hence, the network produces high-dimensional dynamics and low-rank suppression despite the fact that $W$ is low-dimensional and non-normal. The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
$\textrm{col}(U)=\textrm{col}(V)$ $=\textrm{span}\{{\bf{1}}_{e},{\bf{1}}_{i}\}$
$=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}$
where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$. Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24]. As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Figure 16 in Appendix D), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$. As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [44, 45, 46, 47, 48, 49, 50] and widely studied in computational models [21, 40, 24, 3]. Hence, the widely studied theory of balanced network models can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to a particular modular network structure.
Amplification of disordered perturbations in networks with spatial structure
Figure 5: Low-rank suppression in a network with spatial structure. a) Connectivity matrix, $W$. Connection strength is a Gaussian function of distance. b) Singular values of $W$ demonstrate effective low-rank structure. c) A spatially smooth perturbation (red) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (gray) is not. d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a). Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix D). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix D).
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes. Hence, perhaps surprisingly, strongly low-rank networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
Low-rank suppression and high-dimensional dynamics in an epidemiological network.
We next consider a real epidemiological network, specifically a network of high school social contacts [51], which was used in recent theoretical work on low-rank network dynamics [1]. In that work, the authors considered a quenched mean-field reduction of the susceptible-infected-susceptible model,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (11)
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix D). In [1], it was shown that the dynamics generated by Eq. (11) on this network are effectively low-dimensional.
Figure 6: Low-rank suppression and high-dimensional dynamics in a real epidemiological network. a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction. b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
However, Eq. (11) is completely self-contained without any perturbations. Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network produces low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Figure 19b in Appendix D) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
Discussion
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. We also derived conditions under which low-rank networks generate high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics under some assumptions. In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Appendix A for a more complete comparison). Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
Our results also rely on a linearization of the dynamics around a stable fixed point. The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question. However, nonlinear networks can produce any dynamics [29, 30] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks. Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then using bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations. Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d). A large, negative eigenvalue is well-known to produce a suppressive dynamic. However, the results in Figure 2, Figure 12 in Appendix D, and Claim 1 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression. Low-rank suppression occurs in strongly low-rank networks regardless of their eigenvalue spectrum.
Our results have implications for networks in nature and applications. For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. In neuroscience, our results are consistent with the observation that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [16].
Beyond neuroscience, our results imply that perturbations to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space. More generally, perturbations to a low-rank network are more effective when they are not aligned to the network’s low rank structure. This result can be leveraged to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations – except for Figure 6 and Figure 19 in Appendix D – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Figure 19 in Appendix D were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
https://github.com/RobertRosenbaum/HighDimLowDimCode
Acknowledgments
We thank Ashok Litwin-Kumar and David G. Clark for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
Appendix
Appendix A Comparison to previous work on dynamics of low-rank networks
In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 6, 4, 7, 8, 9], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
All of these studies consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model. We next describe the salient properties that distinguish these models from ours.
Thibeault et al. [1] assume that their networks are self-contained and do not receive any time varying external input. Specifically, they explicitly restrict their dynamics to models of the form
$\frac{d{\bf{z}}}{dt}=F({\bf{z}},W{\bf{z}})$
which excludes the possibility of modeling a time-varying external input, like ${\bf{x}}(t)$ in our model. Models of this form cannot describe networks that are part of a larger network, or networks that are modulated by time-varying, external factors. Similar assumptions were made in one study by Landau and Sompolinsky [5]. In other work by Landau and Sompolinsky [3], external input was included, but this input was assumed to be perfectly aligned to the low-rank structure of the connectivity matrix and therefore low-dimensional.
The salient differences between our model and the models considered by Ostojic and colleagues [2, 6, 4, 7, 8, 9] are more subtle. Like us, they consider external inputs that are not aligned to the low-rank part of $W$. Also, like us, they take $W=W_{0}+W_{1}$ where $W_{1}$ is full rank with random entries and the variance of the entries in $W_{1}$ scale like $\mathcal{O}(1/N)$ so that the maximum singular value of $W_{1}$ scales like $\mathcal{O}(1)$.
However, in contrast to our models, the low-rank components of the networks considered by Ostojic and colleagues take the form
$W_{0}=\sum_{\mu=1}^{r}\frac{{\bf{m}}_{\mu}{\bf{n}}_{\mu}^{T}}{N}$ (12)
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (12), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$. This difference in scaling is acknowledged by Ostojic et al. [2, 6, 4, 7, 8, 9] who refer to their networks as “weakly low-rank.”
The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent. For example, recent work by Mastrogiuseppe et al. [10] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values. While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [2], the absence of a random part means that their connectivity matrices are truly low-rank. In that work [10], Mastrogiuseppe et al. showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
For the sake of comparison to the models in [10], let us consider a rank-one network structure like the one in Figure 1. Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
$W_{0}=c{\bf{u}}{\bf{u}}^{T}.$
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 12 in Appendix D which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Figures 7 and 8 in Appendix D).
Figure 7: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure 8: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 7). Similar examples and results are considered in [10]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 8), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 8d; compare to Figure 2 and Figure 12).
Landau and Sompolinsky [5, 3] consider strongly low-rank networks. However, as noted above, their external input, ${\bf{x}}(t)$, is aligned to the low-rank part of $W$ and is therefore low-dimensional itself. Thibeault et al. [1] consider several network models, but their “rank-perturbed Gaussian” model is strongly low-rank and equivalent to the network structure we study in Figure 1. Complicating matters, Thibeault et al. directly compare their rank-perturbed Gaussian model to the networks in Ostojic et al., despite the fact that the scaling of their low-rank parts differ by a magnitude of $\sqrt{N}$. Specifically, in Thibeault et al. [1], the low-rank part is defined by Eq. (12) where $\bf{m}$} and $\bf{n}$} are unbiased Gaussian random vectors. The entries of $\bf{m}$} have $\mathcal{O}(1/N)$ variance while the entries of $\bf{n}$} have $\mathcal{O}(1)$ variance, so the entries of $W_{0}$ have $\mathcal{O}(1/N)$ variance, in contrast to the $\mathcal{O}(1/N^{2})$ variance used by Ostojic et al. [2]. Importantly, this means that the singular values of $W_{0}$ are $\mathcal{O}(\sqrt{N})$ in the rank perturbed Gaussian model analyzed by Thibeault et al. [1], but $\mathcal{O}(1)$ in the models by Ostojic et al. Hence, the rank perturbed Gaussian models considered by Thibeault et al. are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
Appendix B Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–f. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
$\widetilde{C}^{{\bf{x}}}_{jk}(f)=\int_{-\infty}^{\infty}C_{jk}^{{\bf{x}}}(\tau)e^{-2\pi if\tau}d\tau$
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. The assumption of i.i.d. elements is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (6) to get [31, 53, 41, 42, 54]
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}$ (13)
where
$\widetilde{r}(f)=1-2\pi\tau fi$
is scalar.
Eq. (13) quantifies the covariance structure of ${\bf{z}}(t)$ at any given frequency mode, $f$, but we are often specifically interested in the zero-lag temporal covariance,
$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$
From Eq. (13), we therefore have
$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}df$ (14)
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (14) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}\overline{C}^{{\bf{x}}}[I-W]^{-T}.$ (15)
Note that Eq. (15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$. Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate, but our approach here is more mathematically precise.
When principal component analysis is applied to ${\bf{z}}(t)$, the variance explained by each principal component is given by the ordered list of eigenvalues of the covariance matrix, $\overline{C}^{{\bf{z}}}$. From Eq. (15), we see that these eigenvalues are proportional to the squares of the singular values of the matrix
$R=[I-W]^{-1}\sqrt{\overline{C}^{{\bf{x}}}}.$
where $\sqrt{\overline{C}^{{\bf{x}}}}$ is the matrix square root of the symmetric positive definite covariance matrix, $\overline{C}^{{\bf{x}}}$ (distinct from the entry-wise square root in general). Therefore, the decay of the variance explained by each principal component of $\bf{z}$} (as in the blue dots in Figure 1d) are described by the squared singular values of $R$.
In the models we consider, each ${\bf{x}}_{j}(t)$ is an i.i.d. stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$. Therefore, for our models,
$R=[I-W]^{-1}\sqrt{v}$
and therefore
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}[I-W]^{-T}v$
where recall that $v>0$ is a scalar.
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Figure 11), and would not be related to the intrinsic dynamics of interactions within the network.
In conclusion, when ${\bf{x}}(t)$ is high-dimensional, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
$A=[I-W]^{-1}.$
More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{A,k}\,v$
where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values and principal components are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
Therefore, ${\bf{z}}(t)$ is low-dimensional whenever $A$ has a small number of large singular values. As above, we can use the fact that singular values commute with matrix inverses to write our conclusions in terms of
Q=IW.Q=I-W.
Specifically,
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{v}{\sigma^{2}_{Q,N-k}}$ (16)
where $\sigma_{Q,N-k}$ is the (Nk)(N-k)th singular value of $Q$, i.e., the $k$th from the last singular value.
Figure 9: Comparison between theory and simulations for shorter and longer simulations. a) Same as Figure 1d except we added the theoretical values form Eq. (16) (red dots) and the values obtained from Eq. (17) (green dots). The network was simulated for $T=5\times 10^{3}$ time units. b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (16) as red dots (Figure 9a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (15) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (16) would be replaced by
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{U,N-k}}$
where
$U=[I-W]\sqrt{{\overline{C}}^{\bf{x}}}^{-1}$
and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{\hat{R},k}$ (17)
where
$\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (17) gives a much more accurate prediction (Figure 9a, green dots). This confirms that the error in the red dots from Figure 9a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Figure 9b, red dots), further confirming that the errors in Figure 9a are due largely to sampling error.
As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those in Claim 1, we can summarize as follows
Low-rank suppression (as in Figure 1g) occurs whenever Q=IWQ=I-W has at least one asymptotically large singular value.
High-dimensional dynamics (as in Figure 1d) can only occur whenever Q=IWQ=I-W lacks any asymptotically small singular values.
Previously, in proving Claim 1, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
$P=V^{T}U,$
which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ so that $|P_{jk}|=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$. Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$. We next show that if all singular values of $P$ are $\mathcal{O}(1)$ then the network exhibits high-dimensional responses to high-dimensional inputs.
Claim 2.
Under the model and assumptions 1–3 in the main text, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
Proof.
We will prove this claim by proving its contrapositive:
If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2 and Figure 3g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:
If Q=IWQ=I-W has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
We will prove this version of the claim directly. Assume that $Q$ has an asymptotically small singular value. Then there is a $\bf{z}$} satisfying $\|{\bf{z}}\|=1$ and
$(I-W){\bf{z}}=o(1)$
where the notation $o(1)$ means that $\|(I-W){\bf{z}}\|\to 0$ as $N\to\infty$. It is sufficient to show that there is a $\bf{y}$} with $\|{\bf{y}}\|=1+o(1)$ satisfying
$\|P{\bf{y}}\|=o(1).$
We have that
$W_{0}{\bf{z}}+W_{1}{\bf{z}}={\bf{z}}+o(1)$
Multiplying both sides on the left by $U^{T}$ gives
$\Sigma V^{T}{\bf{z}}+U^{T}W_{1}{\bf{z}}=U^{T}{\bf{z}}+o(1)$
Since $W_{1}$ is random and independent from $W_{0}$ and $U$, and since $U^{T}$ projects from $N$ to $r\ll N$ dimensions, the term $U^{T}W_{1}{\bf{z}}$ represents the projection of a random vector onto the low-dimensional column space of $U$, so $\|U^{T}W_{1}{\bf{z}}\|=o(1)$. We therefore have
$\Sigma V^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
We next claim that ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. To see why this is true, note that $UU^{T}{\bf{z}}$ is the orthogonal projection of $\bf{z}$} onto the column space of $W_{0}$. Since $W_{0}$ dominates $W=W_{0}+W_{1}$ (and therefore the the column space of $W$ is dominated by that of $W_{0}$) and $W{\bf{z}}\approx{\bf{z}}$, we may conclude that $\bf{z}$} lies predominantly in the column space of $W$ and therefore of $W_{0}$. In other words, ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. Hence, we can rewrite the equality above as
$\Sigma V^{T}UU^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
which reduces to
$\Sigma P{\bf{y}}={\bf{y}}+o(1).$
where $P=V^{T}U$ and ${\bf{y}}=U^{T}{\bf{z}}$. Note again that $\|{\bf{y}}\|=\|U^{T}{\bf{z}}\|=1+o(1)$ since $\|{\bf{z}}\|=1$ and ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. We then have that
$\|P{\bf{y}}\|=\|\Sigma^{-1}{\bf{y}}\|+o(1)=o(1)$
because $\Sigma^{-1}$ is a diagonal matrix with $o(1)$ terms on the diagonal. ∎
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2 and Figure 3g. Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$.
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (15). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (13) instead of the identity itself, as in Eq. (15). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$). Next, in Appendix C, we generalize away from the assumption of slow inputs using a Laplace transform.
Appendix C The analysis of fast and transient dynamics.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-steady state approximation from Eq. (2) in this case. Instead, we take the Laplace transform in Eq. (6) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (18)
where $\hat{\psi}(s)=\int\psi(t)e^{-st}dt$ is the Laplace transform of $\psi(t)$ and
$\hat{H}(s)$ $=(1+\tau s)^{-1}W$ (19)
$\hat{B}(s)$ $=(1+\tau s)^{-1}.$
Solving the implicit Eq. (18) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-\hat{H}]^{-1}\hat{B}\hat{{\bf{x}}}$. (20)
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$ because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. At mode $s=0$ (the static or “DC” mode), we recover the connectivity from the steady-state or quasi-steady state analysis, $\hat{H}(0)=W$. Similarly, $\hat{B}(0)=1$. Hence, the quasi-steady state analysis from above evaluated how the DC component of the input, $\hat{{\bf{x}}}(0)$, affects the DC component of the dynamics, $\hat{{\bf{z}}}(0)$, even when we do not assume that ${\bf{x}}(t)$ changes slowly. The analysis here is a generalization of the quasi-steady state analysis to non-static modes, $s\neq 0$.
Note also that Eq. (13) from Appendix B can be derived from Eqs. (20) and (19) above by first taking $s=-2\pi if$ in Eq. (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the approach in this section is a generalization of the approach in Appendix B.
To interpret Eq. (20), let us consider two separate limits. At modes that are much slower than the intrinsic timescale of the dynamics, $|s|\ll 1/\tau$, we have that $\hat{H}(s)\approx W$ and $\hat{B}(s)\approx 1$, so
$\hat{{\bf{z}}}(s)\approx[I-W]^{-1}\hat{{\bf{x}}}(s)$
and we recover the results from our quasi-steady state analysis at these modes. In particular, Claim 1 holds and we expect low-rank suppression at mode $s$ whenever $W$ is strongly low rank. While this may at first seem like a repetition of our results from the quasi-steady state analysis, note that we do not need to assume that ${\bf{x}}(t)$ varies slowly here. The Laplace analysis at slow modes ($|s|\ll 1/\tau$) applies even when ${\bf{x}}(t)$ has power at slow and fast modes.
At fast modes, $|s|\gg\tau$, the identity term in Eq. (20) dominates the $\hat{H}$ and $\hat{B}$ terms so that
$\hat{{\bf{z}}}(s)\approx\hat{{\bf{x}}}(s)$
at fast modes. Therefore, fast modes (representing transient responses to transient input dynamics) do not exhibit low-rank suppression, but they do produce high-dimensional responses to high-dimensional stimuli. Modes between these two extremes interpolate between them.
In summary, when our previous assumption that ${\bf{x}}(t)$ varies more slowly than $\tau$ is violated, we expect to see low-rank suppression at slow temporal modes, but not at fast temporal modes.
Additionally, working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (4), which we later linearized. The dynamics in Eq. (4) capture a large class of models, but are not fully general. Specifically, in Eq. (4), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$. However, some systems like integral equations, integro-differential equations, operator equations, and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (4). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (21)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivariant, causal function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (21) encompasses an extremely broad class of systems, including those described by Eq. (4).
We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq. (21). As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$. However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$. We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
A first order Volterra expansion gives Eq. (18) as an implicit equation for ${\bf{z}}(t)$, but the matrices $\hat{H}$ and $\hat{B}$ are no longer defined as above. Instead, the matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (6), we recover the $\hat{H}$ and $\hat{B}$ defined in Eq. (19). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (21).
Appendix D Additional Simulations and Data
This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text.
Figure 10: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure 11: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure 12: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
Figure 13: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$ and the magnitude of $c$ was increased by the same factor (c=40c=-40) to maintain the same degree of low-rank structure (the same ratio between the spectral radius of the random part, $\rho$, and the magnitude of the low-rank part, $|c|$), and the dynamics obey Eq. (9). The instability produced by taking $\rho>1$ generates chaotic dynamics.
Figure 14: Dynamics of an unstable, non-chaotic network with rank-one structure. Same as Figure 1 except $c=10$ and the dynamics obey Eq. (9). The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
Figure 15: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (9), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure 16: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
Figure 17: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
Figure 18: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.